OpenStudy (anonymous):

Solve this quadratic equation. x^2 + 2x - 22 = 0

OpenStudy (anonymous):

What we have to do in this? :/

OpenStudy (anonymous):

I mean solve it by quadratic formula?

OpenStudy (anonymous):

I believe so?

OpenStudy (anonymous):

The answer is x= -1 plus/minus sqrt(23)

OpenStudy (anonymous):

I just need to know how to get there.

OpenStudy (anonymous):

LOL @Euler271 explained you na try to do it yourself :P

OpenStudy (anonymous):

\[x_{1,2} = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] for the equation in the form ax^2 + bx + c

OpenStudy (anonymous):

in the form ax^2 + bx + c = 0 **

OpenStudy (anonymous):

let me know if you want the next step

OpenStudy (anonymous):

LOL let him do it xD

OpenStudy (anonymous):

46 is what I get.

OpenStudy (anonymous):

@tcarroll010 that is confusing :o :/

OpenStudy (anonymous):

a = 1; b = 2; c = -22 \[x_{1,2} = \frac{ -2 \pm \sqrt{2^2 - 4(1)(-22)} }{ 2(1) } = \frac{ -2 \pm \sqrt{92} }{ 2 }\]

OpenStudy (anonymous):

mmhmm

OpenStudy (anonymous):

\[= \frac{ -2 \pm \sqrt{4}\sqrt{23} }{ 2 } = -1 \pm \sqrt{23}\]

OpenStudy (anonymous):

@emcrazy14 Where did the sqrt(4) and 23 come from?

OpenStudy (anonymous):

@who? lol \[\sqrt{a} \times \sqrt{b} = \sqrt{ab}\]

OpenStudy (anonymous):

lol meant @Euler271 anyway, i thought a was 1 and b was 2??

OpenStudy (anonymous):

this is in a general case. for any two numbers a and b

OpenStudy (anonymous):

ah,

OpenStudy (anonymous):

\[\sqrt{92} = \sqrt{4 \times 23} = \sqrt{4} \times \sqrt{23}\]

OpenStudy (anonymous):

Ah. Makes sense. Thanks!!!!!!